# Reconfigurable Four-Photon Interference in a Deployed Metropolitan Fiber Network

### v3

@@ IV.3 Fusion Operation Across the Network
Figure 6: (a) Experimental layout for the four-photon fusion stage locally at UTC. (b) Experimental layout for the four-photon fusion stage across the network. The solid circles in (c)&(d) are the experimental results corresponding to (a)&(b), respectively. A dip in the four-fold coincidence counts indicates multi-photon interference at the fusion BS. (e) Four-fold polarization-resolved coincidence counts over the 16 measurement outcomes, recorded at the dip setting of (d). The solid pink lines in (c)&(d) are Gaussian fits.
With the source and arrival-time synchronization characterized, we demonstrate the four-photon fusion stage described in Section III.A across the network. To highlight network reconfigurability, in this experiment we retain the photons in modes $e$ and $f$ at UTC while routing modes $b$ and $d$ to BQN and TQN, respectively (Fig. 6(b)). We first perform the fusion operation locally by retaining all four photons at UTC (Fig. 6(a)). The local and network fusion results are shown in Figs. 6(c) and 6(d), respectively. As temporal overlap is achieved by adjusting the VDL, multi-photon interference suppresses distinguishable-path coincidence contributions, producing a dip in the measured four-fold coincidence rate. We interpret these dips as evidence of successful temporal overlap and indistinguishability sufficient to observe fusion-stage interference in both local and deployed-fiber configurations.
@@ IV.4 Bell-like Conditional Channel Across the Network
Figure 7: (a) Experimental layout for polarization-controlled generation of the Bell-like conditional channel locally at UTC. (b) Experimental layout for the operation across the network. Data shown in (c)&(d) correspond to (a)&(b), respectively. The solid green and pink circles in (c)&(d) represent the measured four-photon coincidence counts $C_{H_{e}H_{f}}\big|_{V_{b}H_{d}}$ and $C_{H_{e}V_{f}}\big|_{V_{b}H_{d}}$, respectively. The thick solid green and pink curves in panels (c)&(d) denote the theoretical curves for the coincidence probabilities $P_{HH}$ and $P_{HV}$ from Eq. (8), respectively, while the thinner green and pink lines are provided as visual guides for the experimental data.
We next demonstrate polarization-controlled reconfigurable conditional projection corresponding to a Bell-like coincidence channel. We retain the two conditioning photons in modes $b$ and $d$ locally at UTC and route modes $e$ and $f$ through deployed fiber loops to TQN and BQN, respectively, before returning for detection. By projecting the locally retained photons $b$ and $d$ into cross-polarized outcomes (e.g., $V_{b}H_{d}$ or $H_{b}V_{d}$) using the PBSs highlighted in cyan in Figs. 7(a) and 7(b), and by recording four-fold coincidences, we select a conditional subset of remote events whose polarization-correlation dependence follows the theoretical model in Eq. (8). The local and deployed-network configurations are shown in Figs. 7(a) and 7(b), with the corresponding results presented in Figs. 7(c) and 7(d).
@@ IV.5 N00N-like Conditional Channel Across the Network
Figure 8: (a) Experimental layout for polarization-controlled operation of the N00N-like conditional channel locally at UTC. (b) Experimental layout for the operation across the network. Data shown in (c)&(d) corresponds to (a)&(b), respectively. The solid green circles in (c)&(d) represent the measured four-photon coincidence counts with photons $b$ and $d$ projected into the same-polarization modes of $H_{b}H_{d}$. The thick solid pink lines in (c)&(d) are Gaussian fits, while the thinner green lines are provided as visual guides for the experimental data.
@@ V Conclusion and Discussion
In summary, we have demonstrated polarization-controlled reconfigurable four-photon interference and conditional two-photon projection over a metropolitan-scale deployed fiber network using an all-fiber linear-optical platform. By synchronizing photon arrival times across the deployed network and performing polarization-resolved conditioning on locally retained photons, we observe distinct four-fold coincidence interference signatures corresponding to Bell-like and N00N-like conditional channels. These results extend multi-photon interference experiments beyond the laboratory and into a real-world telecommunication infrastructure characterized by substantial loss, polarization drift, and environmental perturbations.
From a systems perspective, this work serves two primary roles. First, it provides a field validation of multi-photon interference as a viable network-level resource. Many advanced quantum-network protocols explicitly rely on the ability to generate and interfere multiple photons with sufficient indistinguishability [10, 11, 12], yet experimental evidence for such processes in deployed fiber has remained limited. Our results show that four-photon interference signatures can be observed on a metropolitan network, while also quantifying the practical limitations imposed by source infidelity, fiber loss, and timing uncertainty.
Second, the polarization-controlled switching between Bell-like and N00N-like conditional coincidence channels demonstrates a programmable network control primitive based entirely on linear optics and local measurements. This capability enables systematic benchmarking of deployed fiber links under different correlation structures and provides a pathway toward adaptive quantum network operation, where conditional multi-photon processes can be selected dynamically in response to network conditions. Notably, these capabilities remain meaningful even in the absence of entanglement verification, making them directly relevant to near-term quantum network development.
@@
In this work, all photons are ultimately detected at UTC after traversing the deployed fiber loops. Extending to true multi-node operation, with coincidence detection at physically separated nodes and independently running clocks, will require tighter timing synchronization (e.g., via the White Rabbit protocol) [25, 26], and a network configuration supporting distributed time-tag correlation without centralized detection. Finally, integrating automatic polarization control protocols (e.g., as in Ref. [27]) will be essential for long-duration stability of multi-photon interference in deployed fiber environments. These developments form the foundation for our next stage of experiments on the EPB-IonQ Bohr-IV quantum network.

### v4

@@ IV.3 Fusion Operation Across the Network
Figure 5: (a) Experimental layout for the four-photon fusion stage locally at UTC. (b) Experimental layout for the four-photon fusion stage across the network. In both (a)&(b), the propagation and detection paths of the four photons are labeled $b$, $d$, $e$, and $f$, consistent with the experimental schematic shown in Fig. 2. In the network setting (b), photons $e$ and $f$ are retained at UTC, while photons $b$ and $d$ are routed through deployed fiber loops to TQN and BQN, respectively, before returning for detection. The solid circles in (c)&(d) are the experimental results corresponding to (a)&(b), respectively. A dip in the four-fold coincidence counts indicates multi-photon interference at the fusion BS. (e) Four-fold polarization-resolved coincidence counts over the 16 measurement outcomes, recorded at the dip setting of (d). The solid pink lines in (c)&(d) are Gaussian fits.
With the source and arrival-time synchronization characterized, we demonstrate the four-photon fusion stage described in Section III.A across the network. To highlight network reconfigurability, in this experiment we retain the photons in modes $e$ and $f$ at UTC while routing modes $b$ and $d$ to BQN and TQN, respectively (Fig. 5(b)). We first perform the fusion operation locally by retaining all four photons at UTC (Fig. 5(a)). The local and network fusion results are shown in Figs. 5(c) and 5(d), respectively. As temporal overlap is achieved by adjusting the VDL, multi-photon interference suppresses distinguishable-path coincidence contributions, producing a dip in the measured four-fold coincidence rate. We interpret these dips as evidence of successful temporal overlap and indistinguishability sufficient to observe fusion-stage interference in both local and deployed-fiber configurations.
@@ IV.4 Bell-like Conditional Channel Across the Network
Figure 6: (a) Experimental layout for polarization-controlled generation of the Bell-like conditional channel locally at UTC. (b) Experimental layout for the operation across the network. In both (a)&(b), the propagation and detection paths of the four photons are labeled $b$, $d$, $e$, and $f$, consistent with the experimental schematic shown in Fig. 2. In the network setting (b), photons $b$ and $d$ are retained at UTC for polarization control, while photons $e$ and $f$ are routed through deployed fiber loops to TQN and BQN, respectively, before returning for detection. Data shown in (c)&(d) correspond to (a)&(b), respectively. The solid green and pink circles in (c)&(d) represent the measured four-photon coincidence counts $C_{H_{e}H_{f}}\big|_{V_{b}H_{d}}$ and $C_{H_{e}V_{f}}\big|_{V_{b}H_{d}}$, respectively. The thick solid green and pink curves in panels (c)&(d) denote the theoretical curves for the coincidence probabilities $P_{HH}$ and $P_{HV}$ from Eq. (13), respectively, while the thinner green and pink lines are provided as visual guides for the experimental data.
We next demonstrate polarization-controlled reconfigurable conditional projection corresponding to a Bell-like coincidence channel. We retain the two conditioning photons in modes $b$ and $d$ locally at UTC and route modes $e$ and $f$ through deployed fiber loops to TQN and BQN, respectively, before returning for detection. By projecting the locally retained photons $b$ and $d$ into cross-polarized outcomes (e.g., $V_{b}H_{d}$ or $H_{b}V_{d}$) using the PBSs highlighted in cyan in Figs. 6(a) and 6(b), and by recording four-fold coincidences, we select a conditional subset of remote events whose polarization-correlation dependence follows the theoretical model in Eq. (13). The local and deployed-network configurations are shown in Figs. 6(a) and 6(b), with the corresponding results presented in Figs. 6(c) and 6(d).
@@ IV.5 N00N-like Conditional Channel Across the Network
Figure 7: (a) Experimental layout for polarization-controlled operation of the N00N-like conditional channel locally at UTC. (b) Experimental layout for the operation across the network. In both (a)&(b), the propagation and detection paths of the four photons are labeled $b$, $d$, $e$, and $f$, consistent with the experimental schematic shown in Fig. 2. In the network setting (b), photons $b$ and $d$ are retained at UTC for polarization control, while photons $e$ and $f$ are routed through deployed fiber loops to TQN and BQN, respectively, before returning for detection. Data shown in (c)&(d) corresponds to (a)&(b), respectively. The solid green circles in (c)&(d) represent the measured four-photon coincidence counts with photons $b$ and $d$ projected into the same-polarization modes of $H_{b}H_{d}$. The thick solid pink lines in (c)&(d) are Gaussian fits, while the thinner green lines are provided as visual guides for the experimental data.
@@
Also note that the purpose of the measurements in Figs. 5(d), 6(d), 7(d) is to demonstrate the presence of the four-photon interference signature under field-deployed network conditions rather than to extract a physical parameter with high statistical precision. The dominant fluctuations arise from slow systematic variations of the deployed fiber network (e.g., polarization drift, fiber stress, temperature variation, and switch routing variation) rather than from Poisson counting statistics. Consequently, statistical error bars derived solely from photon-counting noise would not accurately represent the true experimental uncertainty. Furthermore, the four-photon coincidence rate in the deployed network is extremely low (on the order of $10^{-2}$$10^{-1}\,\mathrm{Hz}$), as shown in Figs. 6(d) and 7(d), acquiring statistically independent repetitions for each data point would require measurement times of several days per interference curve. Instead, the interference visibility is extracted from a global fit to the measured interference pattern, which effectively averages over statistical fluctuations across the dataset.
@@ V Conclusion and Discussion
In summary, we have demonstrated polarization-controlled reconfigurable four-photon interference among three nodes and conditional two-photon projection over a metropolitan-scale deployed fiber network using an all-fiber linear-optical platform. By synchronizing four-photon arrival times across the three distant nodes on the deployed network and performing polarization-resolved conditioning on locally retained photons, we observe distinct four-fold coincidence interference signatures corresponding to Bell-like and N00N-like conditional channels. These results extend multi-photon interference experiments beyond the laboratory and into a real-world, multi-node telecommunication infrastructure characterized by substantial loss, polarization drift, and environmental perturbations.
From a systems perspective, this work serves two primary roles. First, it provides a field validation of multi-photon interference across multiple distant nodes as a viable network-level resource. Many advanced quantum-network protocols explicitly rely on the ability to generate and interfere multiple photons with sufficient indistinguishability [23, 13, 22], yet experimental evidence for such processes in deployed fiber has remained limited. Our results show that four-photon interference signatures can be observed across three distant nodes on a metropolitan network, while also quantifying the practical limitations imposed by source infidelity, fiber loss, and timing uncertainty.
Secondly, the polarization-controlled switching between Bell-like and N00N-like conditional coincidence channels demonstrates a programmable network control primitive based entirely on linear optics and local measurements. This capability enables systematic benchmarking of deployed fiber links under different correlation structures and provides a pathway toward adaptive quantum network operation, where conditional multi-photon processes can be selected dynamically in response to network conditions. Notably, these capabilities remain meaningful even in the absence of entanglement verification, making them directly relevant to near-term quantum network development.
@@
In this work, all photon-arrival times are ultimately detected at UTC after traversing the deployed fiber loops. Extending to nonlocal multi-photon coincidence detection at physically separated nodes with independently running clocks will require tighter timing synchronization (e.g., via the White Rabbit protocol) [19, 16], and a network configuration supporting distributed time-tag correlation without centralized detection. Finally, integrating automatic polarization control protocols (e.g., as in Ref. [4]) will significantly improve long-term stability of multi-photon interference in deployed fiber environments. These developments form the foundation for our next stage of experiments.